# On the Element-Wise Representation and Reasoning in Zero-Shot Image Recognition: A Systematic Survey

### v2

@@ IV-D1 Cross Attention
Cross attention is used to capture similarities between text embeddings and thus correct text representations. BMPNet [109] introduces the cross attention mechanism to improve independent object/state representations. Specifically, for a state, the attention weights of the state and other states are first computed, and then the attention weights of and object are computed, resulting in an enhanced state representation. The object ditto. Unlike object/state relational modeling, CAPE [110] aims to reveal co-occurrence relations between compositions. It adapts the compositional features by calculating the attention weights between the text embeddings of the compositions.
IV-D2 Graph Modeling
@@
As mentioned in the previous section, graphs are excellent at working with structured data and thus can be well adapted to the relationship modeling task. In earlier work, Cruz et al. [16] develop a logic tree to model relations between states and objects. They want the model to learn the notion of connectives such as {and, or, not} to understand the potential connections between nouns. CGE [111] models the entire set of objects and the set of states. It treats each object and state as a graph node and then identifies edges between nodes by the composition labels. Graph convolutional network is later introduced to propagate the information. CVGAE [112] expects the model to be able to determine on its own whether there is a connection between objects and states, and thus be adaptable to the open-world compositional recognition task. It designs a reconstruction task where the reconstruction network recovers the adjacency matrix based on latent variables. GIPCOL [113], on the other hand, fuses prompt learning and graph modeling. It employs CLIP as backbone and takes the graph-encoded text embeddings as prompts to train. The graph neural network is involved in the learning process thereby optimizing the text embeddings to better fit the task.
IV-D3 Contrastive Embedding
Figure 6: A schematic diagram of contrastive embedding [49].
@@ IV-D3 Contrastive Embedding
Contrastive embedding utilizes contrastive learning to optimize the semantic space and implicitly learn the relationships between samples. Contrastive learning [114], as a representative unsupervised learning method, demonstrates powerful semantic representational ability. It aims at clustering similar samples together and separating different samples. The learning mode is to constrain the semantic distance by comparing positive and negative examples, and the learned semantic distance reflects the affinity between the samples, i.e., the semantic relationship. Therefore, many areas exploit contrast learning to optimize semantic embeddings. Unlike the attention mechanism and graph modeling described above, contrast learning does not pre-define the relationship between samples explicitly, but rather iteratively adjusts the semantic distances between samples to place them in appropriate positions. The performance of contrast learning greatly depends on the number and quality of negative examples.
@@
Wei et al. [49] propose a quintuplet loss to optimize the text embedding space. Specifically, they use a sub-network to synthesize state and object texts into compositions. After that, for an image, its label is used as a positive example and supplemented with three negative examples for contrastive learning. The negative examples are text compositions with the same state but different objects, the same object but different states, and different object and state with the target image (Fig. 6). SCEN [115] performs contrastive learning in the visual semantic space. It defines two databases to sample the negative examples, one with the same state as the target image but with different objects, and the other with the same object but with different states. After that it implements contrastive learning in the decoupled state/object embedding space respectively. Yang et al. [116] adopt the similar approach for exploring the relations among states or objects. Meanwhile, they introduce instance-level contrast learning to optimize the compositional embedding space.
IV-E Others
@@
In addition to the above mainstream techniques, many researches provide unique perspectives to address the problem of compositional recognition. TMN [5] presents a gated network that recognizes different compositions by injecting state and object information into the network to determine the switching of pathways. Similarly, TAFE-Net [117] proposes a task-aware learning paradigm that guides the studying process by using task information as cues. Some studies are devoting to analyzing the imbalance issue. MUST [118] argues that the difficulty of recognizing elements varies, e.g., mashed banana is hard to identified as banana but relatively easy to identified as mashed. For this reason, it introduces a weighted loss to balance the effects of different elements. ProLT [119], on the other hand, points out that the data imbalance problem caused by different sample size of elements affects the recognition accuracy. To this end, it similarly proposes a novel weighted loss to solve this problem. CDS-CZSL [120] argues that for an object, the importance of different states is different. For example, sliced is more important than red for strawberry, cause red is a more general state that can be reflected in a wide range of categories. In this regard, it firstly highlights those important states by clustering and contrasting. And then a specificity-refined primitive loss is implemented to optimize the network. Moreover, several studies develop new tasks to address larger challenges [121, 122].
V Foundation model-Based Open-world Recognition
V-A Problem Description
Traditional deep vision models have achieved great results on small closed-form datasets. At the same time, large language models have seen unprecedented growth. There has been a desire to bridge the visual and language barriers and empower models to process both modal data simultaneously. Large pre-trained visual language models (VLMs) offer a promising solution [123]. The motivation is intuitive and succinct, i.e., to learn the matching relations of visual-language in a super-large image-text pair dataset to understand the element-wise knowledge of images. For example, CLIP [29] employs contrastive learning to embed images and text into the same semantic space. Surprisingly, the knowledge learned by VLMs is extremely versatile and transferable, which can be directly applied to the open-world recognition tasks. It is only necessary to associate the text description with the class name of the downstream recognition task, e.g., a photo of a [class], to achieve the purpose of classification and yield excellent results. However, the capabilities of VLMs have not been fully exploited. In recent years, there has been a proliferation of research addressing the generalization capabilities of VLMs in the open world. The mainstream research directions are as follows.
Better VLMs. Upgrading for VLMs has largely been locked into the exploration of pre-training approaches. The motivation for most studies stems from two aspects, data-efficient pre-training [30, 124, 125] and learning-efficient pre-training [126, 127, 128, 129]. The purpose of data-efficient pre-training is to explore efficient training approaches that reduce the data volume. VLMs typically require the support of large-scale datasets, e.g., CLIP trains with 400 million pairs of image-text, which is obviously a time-consuming and labor-intensive project. For this reason, some research has been devoted to developing better learning approaches that maintain excellent representation capabilities while greatly reducing the data volume. The purpose of learning-efficient pre-training can be summarized as learning better representation. Contrastive learning provides a simple yet efficient paradigm for the training of VLMs. However, whether the learned representation is optimal is a long-standing question. Researchers have attempted to parse and tackle it from different perspectives. One is to explore different learning approaches, such as reconstruction [127], generation [128]. Another is to improve the representation space from the perspective of data interaction, e.g., fine-grained image-text interaction [126], data debiasing [129].
Few-Shot Adaptation and OOD Detection. VLMs are widely used in various downstream tasks because of their excellent representation capabilities. Few-shot adaptation refers to the use of VLMs for few-shot recognition tasks [31, 32]. Specifically, by fine-tuning pre-trained VLMs with a small amount of labeled data, a specialized model that performs well on that task can be obtained. The goal of OOD detection, i.e., out-of-distribution detection, is to identify samples that are not in the target categories [130, 131]. Pre-trained VLMs are first fine-tuned on a specific recognition task. In the testing phase, the model needs to identify foreign samples that do not belong to any of the categories and assign them to the unknown class. It is important to note that the OOD does not need to recognize the specific category of the foreign samples.
Zero-Shot Adaptation. Both few-shot adaptation and OOD detection require labeled data as support to fine-tune VLMs, and zero-shot adaptation releases this limitation. Zero-shot adaptation consists of three basic tasks that apply VLMs to open environment recognition tasks under different constraints.
Zero-shot prediction: With only the category names of the task known and without any image data applying VLMs to the downstream task.
Test-Time Adaptation: Know the category names of the task and allow adjustments at the online testing phase.
Unsupervised Adaptation: Know the category names of the task and have part or entire of the unlabeled test data in advance.
In this paper, we will detail the current progress of zero-shot adaptation according to topic.
@@ V-A Problem Description
V-B Zero-Shot Prediction
@@
VLMs have powerful zero-shot prediction capabilities. For a specific task, simply editing the category names into the prompt, e.g., a photo of [class], enables direct classification without fine-tuning. However, there is a performance gap between such a deployment and a model that is fully supervised trained on the downstream task. To address this issue, zero-shot prediction presents promising solutions. A prevalent solution is to design better prompt. The commonly used a photo of [class], while intuitive, is missing a lot of detailed information. For this reason, Menon et al. [33] propose to use a large language model such as GPT-3 [133] to generate multiple word descriptions for classes. In the testing time, the similarity between the image sample and each word description is calculated and the mean value is taken as the discriminative basis. CuPL [134] has a similar idea, which uses this description directly as the prompt for testing. CHiLS [135] also adopts a large language model, which generates multiple subclasses for each category and uses their combined similarity as the discriminative basis. Unlike the design of prompt, SuS-X [136] uses Diffusion [137] to pre-generate multiple pseudo-visual samples for each class, which serves as a support set to assist the judgment at the testing time. CALIP [34] takes the perspective of embedding space, and it introduces the attention mechanism to capture the correlation between text embeddings and visual embeddings, which effectively improves the zero-shot generalization capability.
V-C Test-Time Adaptation
@@
The purpose of Test-Time Adaptation (TTA) is to dynamically adjust the model to progressively adapt to the downstream task during the testing process. The mainstream schemes are categorized into prompt adaptation and cache bank. Prompt adaptation is a technique that combines prompt learning and TTA. During testing, the prompt is set to learnable parameters and the parameters of the model are fixed. By continuously optimizing the prompt representation and thus adapting to the target task. TPT [132] first presents this concept by augmenting the test image into multiple views and then filtering out those with low confidence. The confidence outputs of the remaining augmented samples are averaged as the final prediction and, at the same time, the information entropy is minimized in order to train the prompt (Fig. 7). This paradigm provides a basic template for subsequent research on prompt adaptation. DiffTPT [138] argues that the augmentation approach of TPT lacks diversity, and therefore introduces the diffusion model to generate richer samples for the test image. PromptAlign [139] argues that adjusting the prompt should be done while aligning the distribution of the target and pre-training data. Therefore it introduces additional loss to align the statistics of target and pre-training data, i.e., mean and variance. SwapPrompt [140] and DART [141] use the exponential moving average algorithm to update prompt.
@@
The goal of the Cache bank is to store the tested samples to improve the performance of subsequent tests. An intuitive approach is to cache the previous high-confidence samples as the support set to be queried by subsequent samples. DMN-ZS [142] introduces an attention mechanism to compute the similarity between the new test samples and the support set in order to obtain the weighted features. TDA [143] sets up two cache banks to store the one-hot labels and the negative soft labels of the reliable samples, respectively. The new test samples interact with them to generate predictions. Zhang et al. [144] employ a cache bank to save high-confidence samples to learn high-quality prompts.
In addition, MTA [145] proposes a new loss function to optimize visual features in the semantic space.
V-D Unsupervised Adaptation
@@
Unsupervised adaptation is suitable for offline testing environments where some or all of the test data is available. Although there are no labels available, image samples can provide a priori information to help better prediction. Existing research combines different algorithms to address this problem. UPL [146] fuses prompt learning and self-supervised learning to learn prompt representations through a small number of high-quality pseudo-labels. MUST [147] incorporates image reconstruction techniques to facilitate the model’s understanding of visual information. InMaP [148] introduces prototype learning, which mines a set of visual prototypes from the visual space to replace the text embeddings for prediction. LaFTer [149] trains a text classifier and then uses it to classify images and trains the image encoder with pseudo-labels. ZLaP [150] presents a label propagation algorithm to assign labels to images. In addition, CLIPPR [151] introduces a label distribution prior to predict the data of target domain.
VI Technical Details
@@
[Class/Attr.] means the number of classes and attributes for ZSOR, objects and states for ZSCR, classes for FBOR.
TABLE II: More details on selected advances for zero-shot object recognition and compositional recognition.
@@
Zero-Shot Compositional Recognition Decoupling Nan et al. [93] AAAI 2019 ResNet18 - Contrast UTZ, MIT -
CAUSAL [42] NeurIPS 2020 ResNet18 - Mapping Function UTZ, AOC [Link]
ProtoProp [94] NeurIPS 2021 ResNet18 - Prototype, Graph UTZ, AOC, CGQ [Link]
OADis [98] CVPR 2022 ResNet18 - Attention UTZ, MIT, VAW [Link]
Zhang et al. [101] ECCV 2022 ResNet18 - Attention UTZ, AOC, CLO [Link]
DECA [97] TMM 2022 ResNet18 - Mapping UTZ, MIT [Link]
DRANet [173] ICCV 2023 ResNet18 - Attention UTZ, MIT, CGQ -
ADE [99] CVPR 2023 ViT-B - Attention UTZ, CGQ, CLO [Link]
DFSP [102] CVPR 2023 CLIP - Attention, Prototype UTZ, MIT, CGQ [Link]
Hu et al. [95] AAAI 2023 ResNet18 - Contrast UTZ, CGQ [Link]
CAILA [96] WACV 2024 CLIP - Mapping UTZ, MIT, CGQ, VAW [Link]
Jing et al. [100] AAAI 2024 CLIP - Attention UTZ, MIT, CGQ -
Troika [174] CVPR 2024 CLIP - Attention UTZ, MIT, CGQ [Link]
Dependency Nagarajan et al. [8] ECCV 2018 ResNet18 - Contrast UTZ, MIT [Link]
SymNet [105] CVPR 2020 ResNet18 - Mapping Function UTZ, MIT [Link]
SAD-SP [107] TPAMI 2023 ResNet18 - Mapping Function UTZ, MIT, CGQ -
CoT [103] ICCV 2023 ResNet18, ViT-B - Attention MIT, CGQ, VAW [Link]
CANet [106] CVPR 2023 ResNet18 - Mapping Function UTZ, MIT, CGQ [Link]
Panda et al. [104] PR 2024 ResNet18 - Graph UTZ, MIT, CGQ -
ProCC [108] AAAI 2024 ResNet18 - Mapping Function UTZ, MIT, CGQ [Link]
Relationship Wei et al. [49] ICCV 2019 ResNet18 - Contrast UTZ, MIT -
BMPNet [109] TMM 2021 ResNet18 - Attention UTZ, MIT [Link]
CGE [111] CVPR 2021 ResNet18 - Graph UTZ, MIT, CGQ [Link]
SCEN [115] CVPR 2022 ResNet18 - Contrast UTZ, MIT, CGQ [Link]
CVGAE [112] MM 2022 ResNet18 - Graph UTZ, MIT, CGQ [Link]
CAPE [110] WACV 2023 ResNet18 - Attention UTZ, MIT, CGQ -
Yang et al. [116] TMM 2023 ResNet18 - contrast UTZ, MIT
GIPCOL [113] WACV 2024 CLIP - Graph UTZ, MIT, CGQ [Link]
Others TMN [5] ICCV 2019 ResNet18 - Gated Network UTZ, MIT [Link]
MUST [118] PR 2024 ResNet18 - Weighted Loss UTZ, MIT, CGQ [Link]
ProLT [119] AAAI 2024 ResNet18, CLIP - Weighted Loss UTZ, MIT, CGQ -
CDS-CZSL [120] CVPR 2024 CLIP - Weighted Loss UTZ, MIT, CGQ -
@@
Zero-Shot Object Recognition: [Region] means region-level annotations. [FLO] means Flowers102, [NAB] means NABirds, [DEP] means DeepFashion.
Zero-Shot Compositional Recognition: [UTZ] means UT-Zsppos, [MIT] means MIT-States, [CGQ] means C-GQA, [AOC] means AO-CLEVr, [VAW] means VAW-CZSL, [CLO] means Clothing16K.
TABLE III: More details on selected advances for fundation model-based open-world recognition.

### v3

@@ 4.4.1 Cross Attention
To model the relationship, cross attention is used to capture similarities between text embeddings and thus correct text representations. BMPNet [113] introduces the cross attention mechanism to improve independent object/state representations. Specifically, for a state, the attention weights of the state and other states are first computed, and then the attention weights of an object are computed, resulting in an enhanced state representation. The object ditto. Unlike object/state relational modeling, CAPE [114] aims to reveal co-occurrence relations between compositions. It adapts the compositional features by calculating the attention weights between the text embeddings of the compositions.
4.4.2 Graph Modeling
@@
As mentioned in the previous section, graphs are excellent tools for structured data and thus can be well adapted to the relationship modeling task. In earlier work, Cruz et al. [13] develop a logic tree to model relations between states and objects. They want the model to learn the notion of connectives such as {and, or, not} to understand the potential connections between nouns. CGE [115] models the entire set of objects and the set of states. It treats each object and state as a graph node and then identifies edges between nodes by the composition labels. Graph convolutional network is later introduced to propagate the information. CVGAE [116] expects the model to be able to determine on its own whether there is a connection between objects and states and thus be adaptable to the open-world compositional recognition task. It designs a reconstruction task where the reconstruction network recovers the adjacency matrix based on latent variables. GIPCOL [117], on the other hand, fuses prompt learning and graph modeling. It employs CLIP as the backbone and takes the graph-encoded text embeddings as prompts to train. The graph neural network is involved in the learning process, thereby optimizing the text embeddings to fit the task better.
4.4.3 Contrastive Embedding
@@
Contrastive embedding utilizes contrastive learning to optimize the semantic space and implicitly learn the relationships between samples. Contrastive learning [118], as a representative unsupervised learning method, demonstrates powerful semantic representational ability. It aims at clustering similar samples together and separating different samples. The learning mode is to constrain the semantic distance by comparing positive and negative examples, and the learned semantic distance reflects the affinity between the samples, i.e., the semantic relationship. Therefore, many areas exploit contrast learning to optimize semantic embeddings. Unlike the attention mechanism and graph modeling described above, contrast learning does not pre-define the relationship between samples explicitly but rather iteratively adjusts the semantic distances between samples to place them in appropriate positions. The performance of contrast learning largely depends on the number and quality of negative examples, and therefore, negative sampling is another critical aspect of contrastive learning.
@@
Wei et al. [48] propose a quintuplet loss to optimize the text embedding space. Specifically, they use a sub-network to synthesize state and object texts into compositions. After that, the label of an image is used as a positive example and supplemented with three negative examples for contrastive learning. The negative examples are text compositions with the same state but different objects, the same object but different states, and different objects and states with the target image. SCEN [119] performs contrastive learning in the visual semantic space. It defines two databases to sample the negative examples, one with the same state as the target image but with different objects, and the other with the same object but with different states. After that, contrastive learning is implemented in the decoupled state/object embedding space. Yang et al. [120] adopt a similar approach to explore the relations among states or objects. Meanwhile, they introduce instance-level contrast learning to optimize the compositional embedding space.
4.5 Other Categories
@@
In addition to the above mainstream techniques, many other studies provide unique perspectives to address the problem of compositional recognition. For example, TMN [26] presents a gated network that recognizes different compositions by injecting state and object information into the network to determine the switching of pathways. Similarly, TAFE-Net [121] proposes a task-aware learning paradigm that guides the studying process by using task information as cues. Some methods are devoted to analyzing the imbalance issue. MUST [122] argues that the difficulty of recognizing elements varies, e.g., mashed banana is hard to be identified as banana while relatively easy to be identified as mashed. For this reason, it introduces a weighted loss to balance the effects of different elements. ProLT [123], on the other hand, points out that the data imbalance problem caused by different sample sizes of elements affects the recognition accuracy. To this end, it similarly proposes a novel weighted loss to solve this problem. CDS-CZSL [124] argues that for an object, the importance of different states is different. For example, sliced is more important than red for strawberry, cause red is a more general state that can be reflected in a wide range of categories. In this regard, it first highlights those important states by clustering and contrasting, and then a specificity-refined primitive loss is implemented. Moreover, several studies develop new tasks to address larger challenges [125, 126].
5 Foundation Model-Based Open-World Recognition
5.1 Problem Description
Large Vision-Language Models (VLMs) are advanced deep learning models that seamlessly integrate visual and language processing capabilities [127]. By analyzing extensive datasets of image-text pairs, these models learn the intricate relationships between vision and language, allowing them to understand and generate detailed language descriptions of image content [28]. VLMs are renowned for their robust cross-modal understanding and nuanced element perception, which make them particularly effective for zero-shot adaptation tasks. It is only necessary to associate the text description with the class name of the downstream recognition task, e.g., a photo of a [class], to achieve the purpose of classification and yield excellent results. However, recent studies suggest the capabilities of VLMs have not been fully exploited. Thus the foundation model-based open-world recognition receives more attention.
To provide a deep understanding of this topic, we first compare several prominent VLM-related research tasks. First, Better VLMs explores to improve the pre-training process of VLMs. The motivation for most studies stems from two aspects, i.e., data efficiency and learning efficiency. Data-efficient pre-training aims to reduce the data required while preserving strong representations [29, 128], since large datasets, like CLIP’s 400 million image-text pairs, are resource-heavy. Learning-efficient pre-training is to explore better alternative training methods like reconstruction [129], generation [130]. Next, Few-Shot Adaptation and OOD Detection, in which the former involves fine-tuning pre-trained VLMs with limited labeled data to create specialized models for specific tasks [30, 31], which the latter aims to identify samples not belonging to the target categories [131, 132]. Note that the OOD does not need to recognize the specific category of the foreign samples. Notably, both few-shot adaptation and OOD detection require labeled data to support fine-tuning VLMs. In contrast, Zero-Shot Adaptation releases this limitation. It consists of three basic tasks that apply VLMs to open environment recognition tasks under different constraints, including i) Zero-shot Prediction: with only the category names of the task known and without any image data applying VLMs to the downstream task; ii) Test-time Adaptation: know the category names of the task and allow adjustments at the online testing phase; and iii) Unsupervised Adaptation: Know the category names of the task and have part or entire of the unlabeled test data in advance. In this survey, we refer to the three tasks of zero-shot adaptation, i.e., i), ii) and iii), as the category of foundation model-based open recognition.
@@ 5.1 Problem Description
5.2 Zero-Shot Prediction
@@
VLMs have powerful zero-shot prediction capabilities. For a specific task, simply editing the category names into the prompt, e.g., a photo of [class], enables direct classification without fine-tuning. However, there is a performance gap between such a deployment and a model that is fully supervised and trained on the downstream task. To address this issue, zero-shot prediction presents promising solutions. A prevalent solution is to design better prompts. The commonly used a photo of [class], while intuitive, is missing a lot of detailed information. For this reason, Menon et al. [32] propose to use a large language model such as GPT-3 [134] to generate multiple word descriptions for classes. In the testing time, the similarity between the image sample and each word description is calculated, and the mean value is taken as the discriminative basis. CuPL [135] has a similar idea, which uses this description directly as the prompt for testing. CHiLS [136] also adopts a large language model, which generates multiple subclasses for each category and uses their combined similarity as the discriminative basis. APPLe [137] uses the GPT-3 descriptions to initialize prompt prototypes as the category classifiers. Unlike the design of prompt, SuS-X [138] uses Diffusion [139] to pre-generate multiple pseudo-visual samples for each class, which serves as a support set to assist the judgment at the testing time. CALIP [33] takes the perspective of embedding space, and it introduces the attention mechanism to capture the correlation between text embeddings and visual embeddings, which effectively improves the zero-shot generalization capability.
5.3 Test-Time Adaptation
@@
The purpose of Test-Time Adaptation (TTA) is to dynamically adjust the model to progressively adapt to the downstream task during the testing process. The mainstream schemes are categorized into prompt adaptation and cache bank. Prompt adaptation is a technique that combines prompt learning and TTA. During testing, the prompt is set as learnable parameters and the parameters of the model are fixed. The model continuously optimizes the prompt representation and thus adapts to the target task. TPT [133] first presents this concept by augmenting the test image into multiple views and then filtering out those with low confidence. The confidence outputs of the remaining augmented samples are averaged as the final prediction, and at the same time, the information entropy is minimized to train the prompt (Fig. 6). This paradigm provides a basic template for subsequent research on prompt adaptation. DiffTPT [140] argues that the augmentation approach of TPT lacks diversity and, therefore, introduces the diffusion model to generate richer samples for the test image. PromptAlign [141] argues that adjusting the prompt should be done while aligning the distribution of the target and pre-training data. Therefore it introduces additional loss to align the statistics of target and pre-training data, i.e., mean and variance. SwapPrompt [142] and DART [143] use the exponential moving average algorithm to update the prompt.
@@
The goal of the Cache bank is to store the tested samples to improve the performance of subsequent tests. An intuitive approach is to cache the previous high-confidence samples as the support set to be queried by subsequent samples. DMN-ZS [144] introduces an attention mechanism to compute the similarity between the new test samples and the support set to obtain the weighted features. TDA [145] sets up two cache banks to store the one-hot labels and the negative soft labels of the reliable samples, respectively. The new test samples interact with them to generate predictions. Zhang et al. [146] employ a cache bank to save high-confidence samples from learning high-quality prompts. In addition, MTA [147] proposes a new loss function to optimize visual features in the semantic space.
5.4 Unsupervised Adaptation
@@
Unsupervised adaptation is suitable for offline testing environments where some or all test data is available. Although there are no labels available, image samples can provide a priori information to help better prediction. Existing research combines different algorithms to address this problem. UPL [148] fuses prompt learning and self-supervised learning to learn prompt representations through a small number of high-quality pseudo-labels. MUST [149] incorporates image reconstruction techniques to facilitate the model’s understanding of visual information. InMaP [150] introduces prototype learning, which mines a set of visual prototypes from the visual space to replace the text embeddings for prediction. LaFTer [151] trains a text classifier and then uses it to classify images and trains the image encoder with pseudo-labels. ZLaP [152] presents a label propagation algorithm to assign labels to images. In addition, CLIPPR [153] introduces a label distribution prior to predicting the data of the target domain.
6 Technical Details
@@
[Class/Attr.] means the number of classes and attributes for ZSOR, objects and states for ZSCR, classes for FBOR. Click the [link] to check more information in the corresponding repository.
TABLE II: More details on selected advances for zero-shot object recognition.
@@
[Region] means region-level annotations. [FLO] means Flowers102, [NAB] means NABirds, [DEP] means DeepFashion. Click the [link] to check more information in the corresponding repository.
TABLE III: More details on selected advances for zero-shot compositional recognition.
Topic Tax. Name Venue Year Backbone KeyWords Dataset Code
Zero-Shot Compositional Recognition Decoupling Nan et al. [96] AAAI 2019 ResNet18 Contrast UTZ, MIT -
CAUSAL [41] NeurIPS 2020 ResNet18 Mapping Function UTZ, AOC [Link]
ProtoProp [97] NeurIPS 2021 ResNet18 Prototype, Graph UTZ, AOC, CGQ [Link]
OADis [101] CVPR 2022 ResNet18 Attention UTZ, MIT, VAW [Link]
Zhang et al. [104] ECCV 2022 ResNet18 Attention UTZ, AOC, CLO [Link]
DECA [100] TMM 2022 ResNet18 Mapping UTZ, MIT [Link]
DRANet [175] ICCV 2023 ResNet18 Attention UTZ, MIT, CGQ -
ADE [102] CVPR 2023 ViT-B Attention UTZ, CGQ, CLO [Link]
DFSP [105] CVPR 2023 CLIP Attention, Prototype UTZ, MIT, CGQ [Link]
Hu et al. [98] AAAI 2023 ResNet18 Contrast UTZ, CGQ [Link]
CAILA [99] WACV 2024 CLIP Mapping UTZ, MIT, CGQ, VAW [Link]
Jing et al. [103] AAAI 2024 CLIP Attention UTZ, MIT, CGQ -
Troika [106] CVPR 2024 CLIP Attention UTZ, MIT, CGQ [Link]
Dependency Nagarajan et al. [6] ECCV 2018 ResNet18 Contrast UTZ, MIT [Link]
SymNet [109] CVPR 2020 ResNet18 Mapping Function UTZ, MIT [Link]
SAD-SP [111] TPAMI 2023 ResNet18 Mapping Function UTZ, MIT, CGQ -
CoT [107] ICCV 2023 ResNet18, ViT-B Attention MIT, CGQ, VAW [Link]
CANet [110] CVPR 2023 ResNet18 Mapping Function UTZ, MIT, CGQ [Link]
Panda et al. [108] PR 2024 ResNet18 Graph UTZ, MIT, CGQ -
ProCC [112] AAAI 2024 ResNet18 Mapping Function UTZ, MIT, CGQ [Link]
Relationship Wei et al. [48] ICCV 2019 ResNet18 Contrast UTZ, MIT -
BMPNet [113] TMM 2021 ResNet18 Attention UTZ, MIT [Link]
CGE [115] CVPR 2021 ResNet18 Graph UTZ, MIT, CGQ [Link]
SCEN [119] CVPR 2022 ResNet18 Contrast UTZ, MIT, CGQ [Link]
CVGAE [116] MM 2022 ResNet18 Graph UTZ, MIT, CGQ [Link]
CAPE [114] WACV 2023 ResNet18 Attention UTZ, MIT, CGQ -
Yang et al. [120] TMM 2023 ResNet18 contrast UTZ, MIT
GIPCOL [117] WACV 2024 CLIP Graph UTZ, MIT, CGQ [Link]
Others TMN [26] ICCV 2019 ResNet18 Gated Network UTZ, MIT [Link]
MUST [122] PR 2024 ResNet18 Weighted Loss UTZ, MIT, CGQ [Link]
ProLT [123] AAAI 2024 ResNet18, CLIP Weighted Loss UTZ, MIT, CGQ -
CDS-CZSL [124] CVPR 2024 CLIP Weighted Loss UTZ, MIT, CGQ -
[UTZ] means UT-Zsppos, [MIT] means MIT-States, [CGQ] means C-GQA, [AOC] means AO-CLEVr, [VAW] means VAW-CZSL, [CLO] means Clothing16K. Click the [link] to check more information in the corresponding repository.
6.1 Standardized Datasets
In Table I, we list the commonly used standardized datasets for three tasks ZSOR, ZSCR, and FBOR, respectively. Specifically, there are 8 commonly used datasets used for ZSOR, including fine-grained datasets such as CUB [7], which is a bird dataset; SUN [8], which is a scene dataset, and coarse-grained datasets such as AWA [156], which is a dataset of various types of animals. Most of the datasets contain detailed attribute annotations. Note that Flowers102 [154] and NABirds [155] do not have attribute annotations. Flowers102 contains textual descriptions, i.e., each image is described by several sentences. NABirds has region annotations, which mark the parts of the bird and are typically used in visual component analysis methods. For ZSCR, there are 6 commonly used datasets. Different from ZSOR, the column [Class/Attr.] indicates the number of objects and the number of states they contain. It can be seen that their compositional sizes cover different orders of magnitude, ranging from $3/8$ (AO-CLEVr [41]), which means up to 24 compositions, to $674/413$ (C-GQA [115]), which means 278,362 compositions. Last, we show 15 commonly used datasets for FBOR that cover different categories and different data sizes. ImageNet [1] includes data with up to 1000 categories, while EuroSAT [169] and CIFAR-10 [170] have only 10 categories. The data types are also diverse, including food [164], flowers [154], pets [167], transportation [165, 166], etc. They are also fine-grained datasets. In addition, DTD [168] is a texture dataset, and EuroSAT [169] is a remote sensing dataset.
6.2 Implementation
To facilitate the deep understanding and fast reusing of ZSIR approaches, this survey further provides the implementation details of most representative methods in Tables II, III, and IV. Specifically, Table II and III show the implementation details of the related studies of ZSOR and ZSCR, including the backbone, auxiliary, dataset, and code addresses. It can be seen that commonly used backbones include ResNet [176], VGG [177], GoogleNet [178], AlexNet [179], ViT [69], and CLIP [28]. [Auxiliary] refers to additional information such as region-level annotations. We provide the dataset and code address to facilitate access to the performance of the method as well as its implementation. Finally, we label the study with up to two keywords based on the specific technique used.
In Table IV, we show more details of the related work in foundation model-based open-world recognition. We list their visual and text encoders, as well as auxiliary information. GPT-3 [134] and Diffusion [139] are the exotic models. [Statistic] denotes the mean and variance of the pre-training data. [Distribution] denotes the prior distribution of labels. We also provide their code repositories for quick reference.
TABLE IV: More details on selected advances for fundation model-based open-world recognition.